Nick went on a bike ride to a store 5 miles away. It took him half an hour to get there and 2/5 of an hour to get back. what was nick's average rate of speed(miles per hour) for the entire trip? answer in decimal form and round to the nearest hundredth when necessary.
How do you do this problem?

Answer by TimothyLamb(4379)

s = d/t
nick avg speed = s = 10/(1/5 + 2/5)
s = 10/(3/5)
s = (10*5)/3
s = 50/3
s = 16.6666666
---
answer:
s = 16.67 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php